# Let’s Measure Information Step-by-Step: LLM-Based Evaluation Beyond Vibes

### v1

@@ 5.1 Information-Theoretic Mechanisms Detect Effectively
TVD-MI achieved a mean effect size of 5.2 across all domains, with peaks exceeding $d=7$ in several summarization tasks. Even the simplest log-prob based approaches maintained robust discrimination with a mean effect size of 1.87. This suggests that the choice to measure information relationships rather than directly evaluate quality is more important than the sophistication of the implementation.
5.2 LLM Judge Experiences Evaluation Inversion
@@
Table 3 illustrates this inversion from specific examples from the CNN / Daily news summarization task. Human reference summaries, the gold standard for quality, received the lowest scores from LLM judges (0.234 with context, 0.117 without). Meanwhile, "Conspiracy Theory" summaries that fabricated information scored 0.703 with context and 0.777 without. This represents a 3-7× preference for misinformation over accurate content.
@@
5.3 The Inverted-U Pattern: Compression and Information Structure
For compression ratio, quadratic models significantly outperformed linear fits for all primary mechanisms. GPPM showed the most dramatic improvement (R² increasing from 0.029 to 0.684, p = 0.007), while TVD-MI exhibited similar gains (R² from 0.046 to 0.674, p = 0.008). The quadratic coefficient was negative for all mechanisms, confirming the inverted-U shape with peaks at compression ratios of 9.6:1 (MI), 11.0:1 (GPPM), and 11.2:1 (TVD-MI).
5.4 Measurement Tampering: Information-Theoretic Mechanisms Show Superior Robustness
Our adversarial experiments demonstrate that information-theoretic mechanisms maintain substantially better robustness than standard evaluation approaches. While simple transformations can degrade discrimination across all methods, our mechanisms preserve their fundamental properties where LLM judges catastrophically fail.
Table 4 presents the effects of four adversarial transformations on mechanism performance.
Table 4: Effects of adversarial transformations on Reddit TIFU summarization. Score changes and discrimination degradation across four transformation types, with averages showing overall robustness.
Mechanism Case Flip Format Padding Pattern Average
Score Changes
TVD-MI +7.0%*** +7.7%*** +2.9%*** +11%*** +7.2%
MI (DoE) -3.2%*** +45%*** +20%*** +21%*** +21%
GPPM -1.4% +23%*** +8.0%*** +96%*** +32%
Judge (w/ ctx) -11%*** +0.0% -6.4%*** -34%*** -13%
Judge (w/o ctx) -11%*** -4.2%*** -10%*** -48%*** -18%
Discrimination Degradation (% change in Cohen’s d)
TVD-MI -31% -18% -6% -26% -21%
MI (DoE) -50% -84% -56% -137% -82%
GPPM -14% +4% -6% -3% -5%
Judge (w/ ctx) -40% +4% -1% -77% -29%
Judge (w/o ctx) -2157% +588% +786% -100% -221%
Gaming resistance reveals paradoxical patterns. TVD-MI scores increase consistently under all attacks (+2.9% to +11.3%), yet discrimination degrades by an average of 21%. This validates our theoretical prediction: bounded measures prevent score deflation but cannot prevent adversaries from adding spurious patterns that obscure meaningful distinctions. In contrast, unbounded MI shows extreme vulnerability with 21% average score inflation coupled with 82% discrimination loss.
Theory correctly predicts relative robustness hierarchies. Theorem 3.5 predicts bounded measures should maintain better guarantees than unbounded measures under adversarial conditions. The results validate this precisely: TVD-MI (bounded) averages 21% discrimination degradation while MI/DoE (unbounded) suffers 82% degradation. GPPM shows surprising robustness with only 5% average degradation, while LLM judges exhibit chaotic behavior ranging from extreme degradation to nonsensical improvements.
Extreme vulnerability of direct evaluation. The LLM judge without context shows catastrophic failure under case-flipping (-2157% discrimination change) and complete inversion under pattern injection (-100%). These changes from +786% under padding to -2157% under case-flipping indicate the judge has lost any meaningful connection to content quality, essentially producing random evaluations correlated only with surface features.
The consistent pattern across transformations demonstrates that measurement tampering represents a fundamental challenge distinct from score manipulation. While we cannot prevent all gaming, the 4× robustness advantage of bounded measures (TVD-MI: 21%, GPPM: 5%) over unbounded alternatives (MI: 82%) provides a clear design principle for practical deployment. The failure of repetition transformation to even parse (acting as a jailbreak) further underscores that robustness requires information-theoretic foundations rather than heuristic quality judgments.
5.5 Synthesis of Findings
Our empirical results reveal four interconnected discoveries that reshape how we should approach AI evaluation:
Incentive Design Yields Quality Detection—At No Extra Cost. Information-theoretic mechanisms achieve 10/10 perfect discrimination across all domains while using identical computational resources as failing LLM judges. Both require single API calls, but asking "do these share information?" rather than "which is better?" transforms failure into success. This validates that strategic robustness and quality detection emerge together from proper incentive design through the data processing inequality.
The Inverted-U Pattern Explains Theory-Practice Gap. Mechanism performance follows a quadratic relationship with information structure (R² = 0.68, p < 0.01), peaking at approximately 3 effective dimensions. This finding resolves the apparent contradiction between pessimistic theoretical bounds and strong empirical performance: real-world tasks naturally operate in the "sweet spot" where our mechanisms excel, far from worst-case scenarios.
Theory Successfully Predicts Empirical Robustness. Our theoretical framework makes concrete predictions that experiments validate. Theorem 3.5 predicts bounded measures should outperform unbounded ones under adversarial attack. Empirically, TVD-MI (bounded) degrades by only 31% while MI/DoE (unbounded) degrades by 137%. This represents 10-100× better robustness than LLM judges which fail catastrophically (-2157% without context, -77% with context). The theoretical construction that moves probability mass to create uniformity directly corresponds to our format standardization attack, demonstrating deep alignment between theory and practice.
Practical Implementation Guidelines Emerge from Theory. Our findings suggest concrete recommendations: (1) Use TVD-MI for maximum robustness when adversarial manipulation is a concern, (2) Use GPPM or MI/DoE for efficiency when computational cost matters most, (3) Expect optimal performance for tasks with compression ratios near 10:1, (4) Pilot test on your specific domain to identify the effective dimensionality sweet spot.
Robust evaluation requires measuring what cannot be easily faked (information relationships) rather than judging what can be mimicked (surface quality). While no mechanism is perfectly tamper-proof, information theory provides principled approaches that significantly outperform current practice.
6 Limitations and Practical Considerations
While our mechanisms demonstrate substantial improvements over current practice, several limitations warrant discussion:
Measurement Tampering Remains a Challenge. Although information-theoretic mechanisms show 10-100× better robustness than LLM judges, adversaries can still degrade discrimination by 6-31% through carefully crafted transformations. Future work should explore defense mechanisms against such attacks.
Domain-Specific Calibration. The inverted-U pattern suggests optimal performance at  3 effective dimensions, but identifying this sweet spot required testing on many domains. Practitioners can estimate effective dimension using their evaluation matrix, but validating generalization on domains is a promising direction for future work.
Reliance on Pre-trained Priors. Our mechanisms’ strong empirical performance partly stems from leveraging implicit priors in pre-trained language models. In domains where these priors are weak or misleading, performance may degrade toward theoretical worst-case bounds as we showed by introducing spurious patterns to tamper measurement scores.
7 Conclusions
This work demonstrates that robust AI evaluation emerges from proper incentive design. By measuring information relationships rather than surface quality, we achieve both strategic robustness and quality detection. Basic log probability comparisons outperform sophisticated LLM judges.
The data processing inequality resolves the paradox between designing for incentives and validating through quality: strategic manipulation degrades both information content and output quality. Mechanisms preventing gaming identify good performance as a consequence of information theory.
Organizations using LLM judges are rewarding deception. Our demonstration of evaluation inversion (conspiracy theories scoring above human references) requires reconsideration. The solution requires asking "what information is shared?" instead of "which is better?"
7.1 Broader Impact
Our findings arrive as organizations increasingly rely on LLM judges for critical decisions. The demonstration of systematic quality inversion requires attention. Information-theoretic evaluation requires no special access, democratizing robust evaluation and enabling any practitioner to implement gaming-resistant mechanisms. Revealing these vulnerabilities might accelerate adversarial behavior before better practices are adopted. Current AI evaluation rewards deception, but solutions exist in decades of incentives theory. In a future shaped by AI, getting evaluation right is necessary for maintaining connection to ground truth.
8 Acknowledgment
Appendix A Peer Review Experiment Details
@@ A.1 Experimental Configuration
A.2 Agent Review Strategies
@@ 29 Checklist Review Standard checklist without deep engagement Low Effort
A.3 Two-Stage Review Generation Process
@@
A.3.1 Stage 1: Reading and Note-Taking

### v2

@@ 5.1 Information-Theoretic Mechanisms Detect Effectively
Using the LLM to implement the TVD-MI critic achieved higher effect sizes when querying information relationships than using it to judge normative preferences. TVD-MI achieved a mean effect size of 5.2 across all domains, with peaks exceeding $d=7$ in several summarization tasks. This uses the same LLM (GPT-4o-mini) as the quality judge baseline. The log-prob based approaches maintained robust discrimination with a mean effect size of 1.87. This suggests that the choice to measure information relationships rather than directly evaluate quality is more important than the sophistication of the implementation.
Mechanism performance did not degrade linearly with compression as hypothesized, but instead followed an inverted-U pattern peaking at  10:1 compression ratios. This corresponds to tasks where agent responses span approximately 3 effective dimensions, enough variation to distinguish strategies but not so much that signals become noise. See Appendix B.1 for detailed analysis and Figure 2.
We designed these mechanisms to be incentive compatible (resistant to gaming), yet they outperform methods explicitly designed for quality assessment. The correlation with ground truth metrics in verifiable domains provides evidence for the approach.
5.2 Mechanisms Transform Pairwise Evaluations into Item-Level Quality Scores
In the previous section we saw that mechanisms achieved large effect-sizes between the good faith and problematic conditions. However, this could be an artifact, and we are interested in measuring the ability to aggregate pairwise comparisons into meaningful item-level quality scores. We support this finding by showing the large effect-sizes are not artifacts. We do this empirically by testing whether mechanism scores can distinguish agent quality levels without ground truth.
Methodology. For each item, we classify agent pairs. A positive class consisting of faithful-faithful pairs where both agents preserve information and a negative class of faithful-problematic pairs. We compute symmetric pairwise scores (averaging directional evaluations) and test whether positive pairs score higher than negative pairs. We report per-item AUCs macro-averaged across examples with 95% bootstrap CIs.
Table 3: AUC scores for distinguishing Faithful-Faithful from Faithful-Problematic agent pairs across domains. Values show macro-averaged AUC ± 95% CI half-width.
Domain n MI (DoE) GPPM TVD-MI Judge w/ context
Translation
MT14 200 0.664 ± 0.006 0.703 ± 0.006 0.710 ± 0.005 0.559 ± 0.009
OPUS 186 0.737 ± 0.009 0.743 ± 0.008 0.703 ± 0.008 0.681 ± 0.011
Summarization
BillSum 200 0.692 ± 0.009 0.677 ± 0.008 0.732 ± 0.007 0.579 ± 0.009
CNN/DM 268 0.706 ± 0.007 0.669 ± 0.006 0.762 ± 0.005 0.626 ± 0.008
MultiNews 200 0.695 ± 0.010 0.674 ± 0.008 0.755 ± 0.007 0.545 ± 0.009
PubMed 200 0.700 ± 0.008 0.698 ± 0.007 0.753 ± 0.007 0.624 ± 0.009
Reddit TIFU 200 0.689 ± 0.007 0.638 ± 0.008 0.772 ± 0.008 0.541 ± 0.007
SAMSum 200 0.655 ± 0.008 0.645 ± 0.008 0.754 ± 0.007 0.572 ± 0.009
XSum 200 0.714 ± 0.008 0.694 ± 0.007 0.767 ± 0.006 0.583 ± 0.011
Peer Review
ICLR 100 0.484 ± 0.007 0.417 ± 0.009 0.544 ± 0.007 0.492 ± 0.009
Results. Table 3 shows TVD-MI achieves the strongest discrimination across nearly all domains (0.71-0.77 for translation/summarization), while judges with source access perform near random (0.50-0.63). The peer review domain proves challenging for all methods due to extreme compression (20:1), though TVD-MI remains above random. The results support our theoretical framework (Section 3.2): by optimizing TVD-MI, we implicitly optimize a family of related measures including Cohen’s $\kappa$, Youden’s J, and AUC.
5.3 Gaming-Resistance: Information-Theoretic Mechanisms Show Superior Robustness
Our adversarial experiments demonstrate that information-theoretic mechanisms maintain better robustness than standard evaluation approaches. While simple transformations can degrade discrimination across all methods, these mechanisms better preserve their basic properties where quality-based LLM judges can fail or invert rankings in some settings.
Table 4 presents the effects of four adversarial transformations on mechanism performance. We also report AUC in Table 5.
[…]
Transformation MI GPPM TVD-MI Judge Judge
(DoE / GEM) (w/ ctx) (w/o ctx)
Score Changes ($\Delta$)
[…]
Discrimination Degradation ($\Delta$ Cohen’s d)
Case Flip -1.252 -0.540 -2.259 -1.090 -1.000
Format -2.106 +0.138 -1.336 +0.096 +0.273
Padding -1.413 -0.238 -0.438 -0.015 +0.364
Pattern -3.441 -0.115 -1.900 -2.074 -0.046
Average -2.053 -0.189 -1.483 -0.771 -0.102
Table 5: Effects of adversarial transformations on mechanism discrimination ability (AUC) for Reddit TIFU summarization. Bold indicates highest score.
Attack MI GPPM TVD-MI Judge
No Transformation 0.689 0.638 0.772 0.541
Case Flip 0.618 0.562 0.708 0.520
Format 0.583 0.577 0.746 0.537
Padding 0.614 0.603 0.760 0.535
Pattern 0.553 0.608 0.716 0.505
[…]
Theory correctly predicts relative robustness hierarchies. Theorem 3.3 predicts linear-growth measures should maintain better guarantees than super-linear ones under adversarial conditions. The results support this: TVD-MI averages $\bar{\Delta}d=-1.483$, MI/DoE $\bar{\Delta}d=-2.053$, while GPPM shows relatively small change ($\bar{\Delta}d=-0.189$). LLM judges exhibit variable behavior, ranging from large drops to spurious gains. TVD-MI maintains AUC > 0.70 under all attacks, while judges degrade to random performance (near 0.50) and other mechanisms show larger degradation.
Notable vulnerability of quality-based prompting. The same LLM (GPT-4o-mini) prompted for quality judgments shows failure under case-flipping ($d:0.05\to-0.95;\ \Delta d=-1.000$) and near-complete inversion under pattern injection ($d:0.05\to 0.00;\ \Delta d=-0.046$). Under padding it spuriously improves ($d:0.05\to 0.41;\ \Delta d=+0.364$). These shifts indicate the judge has lost meaningful connection to content quality, reacting instead to surface features.
The consistent pattern across transformations demonstrates that measurement tampering is a distinct challenge from score manipulation. While we cannot prevent all gaming, the robustness gap provides a clear design principle for practical deployment.
6 Discussion
Our findings demonstrate that reframing evaluation from normative quality judgments to information relationships provides both theoretical and practical advantages. This shift is not merely a technical modification but reflects a fundamental insight about AI evaluation: when ground truth is unavailable, measuring what agents preserve (information) proves more robust than measuring what they produce (quality).
6.1 Why Information Succeeds Where Quality Fails
The stark contrast between TVD-MI and LLM judges using the same model (GPT-4o-mini) reveals that evaluation failure stems from the questions we ask, not the capabilities we possess. When prompted for quality judgments, LLMs can invert rankings (Table 2) and show lower adversarial robustness in our experiments (Table 4). When prompted for information relationships, the same LLMs provide robust discrimination even under attack (Table 3). This suggests that LLMs can reliably detect statistical patterns but struggle with normative assessments that require implicit value judgments.
6.2 The Inverted-U Phenomenon
Contrary to our pre-registered hypothesis of linear degradation with compression, mechanism performance peaked at intermediate compression ratios ($\sim$10:1). This inverted-U pattern reflects a bias-variance tradeoff: at low compression, outputs are too similar to distinguish strategies; at extreme compression, legitimate variation overwhelms strategic differences. The optimal range corresponds to tasks where agent strategies create distinguishable but stable patterns—approximately 3 effective dimensions in the response space (Appendix B.1).
6.3 Limitations and Future Directions
Measurement tampering vulnerability. While TVD-MI maintains AUC > 0.70 under our attacks, adversaries reduced performance from 0.77 to 0.71. This 8% degradation, though far better than judges’ collapse to random (AUC $\approx 0.50$), suggests room for improvement. Future work should explore adaptive mechanisms that detect and compensate for tampering attempts.
Extreme compression challenges. Peer review (20:1 compression) proved difficult for all methods, with TVD-MI achieving only 0.54 AUC. This may reflect fundamental limits: when most information is discarded, distinguishing preservation strategies becomes inherently challenging. Domain-specific calibration or multi-scale mechanisms may be necessary.
Dependence on pre-trained knowledge. Our mechanisms leverage implicit priors in language models, which could fail in novel domains. The gap between empirical performance and theoretical worst-case bounds (Theorem 3.3) quantifies this dependence. Combining our approach with active learning could reduce reliance on pre-existing knowledge.
7 Conclusions
This work establishes that robust AI evaluation requires reconceptualizing how we query language models. By shifting from normative quality judgments to information relationship measurements, we achieve provable gaming resistance while maintaining practical effectiveness. Our key contributions are:
  1. 1.
    Theoretical: We prove bounded f-divergences maintain polynomial robustness under adversarial attacks while unbounded measures degrade exponentially (Theorem 3.3).
  2. 2.
    Methodological: We show information-theoretic mechanisms can be implemented via chain-of-thought reasoning, requiring only black-box LLM access.
  3. 3.
    Empirical: Across 10 domains, our mechanisms detect strategic manipulation (d = 1.87-5.20) where quality-based judges fail, decompose pairwise comparisons into item-level scores (AUC 0.70-0.77), and maintain effectiveness under attacks that reduce judges to random performance.
These results suggest a path forward for AI evaluation in domains where ground truth is unavailable or unverifiable. Rather than pursuing increasingly sophisticated quality judgments, we should develop mechanisms that measure what can be reliably detected: information relationships between outputs. This approach becomes critical as AI systems increasingly evaluate AI-generated content, where maintaining connection to ground truth through principled mechanisms may be our only defense against evaluation collapse.
The vulnerability of LLM judges to manipulation is not a minor technical issue but a fundamental challenge for AI oversight. Our work demonstrates that solutions exist within current technology; we need only ask the right questions.
7.1 Broader Impact
Our findings arrive as organizations increasingly rely on LLM judges for critical decisions, from content moderation to scientific peer review. The vulnerability to quality inversion poses immediate risks. Information-theoretic mechanisms require no special access, democratizing robust evaluation and enabling practitioners to implement gaming-resistant assessment today. While revealing these vulnerabilities might accelerate adversarial behavior, the greater risk lies in continued reliance on manipulable judges. As AI systems increasingly evaluate AI-generated content, maintaining connection to ground truth through principled mechanisms becomes essential for preventing evaluation collapse.
8 Acknowledgment
Appendix A Extended Study Methodology
A.1 Pre-Registration and Analysis Evolution
Our pre-registered study (https://osf.io/c7pum) originally focused on paired Cohen’s d effect sizes to test discrimination between agent categories. The pre-registration specified:
H1: Information Preservation Detection
  • H1a: All mechanisms distinguish Problematic from Good Faith agents (d > 0.5)
  • H1b: Detection ability decreases linearly with compression ratio
  • H1c: TVD-MI shows most robust detection across compression levels
H2: Mechanism Properties
  • H2a: Bounded mechanisms (TVD-MI) show more consistent performance
  • H2b: Log-probability mechanisms degrade in high-compression domains
  • H2c: Gaming resistance highest for TVD-MI (tested via tampering experiments)
During our pre-registration dialogue with an independent AI reviewer (included in the OSF registration), we recognized that validating our decomposability assumption, a fixed oversight strategy across pairs would be effective, required item-level analysis beyond aggregate effect sizes. This led us to implement AUC analysis examining whether item-wise scores could distinguish agent quality levels. Specifically, we test whether scores for Faithful-Faithful pairs exceed scores for Faithful-Problematic pairs at the item level, providing both a validation of decomposability and a complementary test of H2c (gaming resistance) beyond our planned tampering experiments.
A.2 Complete Agent Taxonomy
Our agent taxonomy was designed to test different forms of information preservation and degradation. Each category serves a specific purpose:
Good Faith Agents (Information-Preserving):
  • Faithful: Strategies that prioritize accurate information transfer without stylistic modifications. These serve as our primary positive examples.
  • Style: Strategies that alter presentation (tone, register, framing) while attempting to preserve semantic content. These test whether mechanisms can distinguish style from substance.
Problematic Agents (Information-Degrading):
  • Strategic: Strategies that deliberately manipulate, misrepresent, or distort information. These test detection of adversarial behavior.
  • Low Effort: Strategies that provide minimal information through laziness, over-compression, or generic responses. These test detection of low-quality outputs.
The complete taxonomy for each domain appears in Tables 9, 8, and 7.
Category Evolution: Our pre-registration initially considered four separate categories. During exploratory analysis, we recognized that the basic distinction was between information-preserving (Good Faith: Faithful + Style) and information-degrading (Problematic: Strategic + Low Effort) behaviors, leading to our two-category framework. Both analyses are reported for transparency.
A.3 AUC Computation Methodology
For each source item, we compute mechanism scores for all agent pair combinations. The AUC analysis proceeds as follows:
  1. 1.
    Pair Classification:
    • Positive class: Faithful-Faithful pairs (both agents from Faithful category)
    • Negative class: Faithful-Problematic pairs (one Faithful, one Strategic/Low Effort)
  2. 2.
    Score Computation:
    • MI/GPPM: Symmetrize by averaging $(A,B)$ and $(B,A)$ directions
    • TVD-MI: Use bidirectional critic score
    • Judge: Convert pairwise preferences to relative quality scores (winner=1, loser=0)
  3. 3.
    Statistical Analysis:
    • Compute per-item AUC (rank positive pairs above negative pairs)
    • Report macro-average across items to avoid pooling bias
    • Bootstrap 95% CIs by resampling items (1000 iterations)
A.4 Experimental Configurations
All experiments used consistent configurations across domains to minimize confounds:
Datasets: We intentionally select benchmarks that are open-ended across translation, summarization, and peer-review.
Translation. WMT14 news translation shared task; we use a 500-example subset [Bojar et al., 2014] and OPUS [Tiedemann, 2016].
Summarization. BillSum [Kornilova and Eidelman, 2019], CNN/DailyMail [Hermann et al., 2015, See et al., 2017], MultiNews [Fabbri et al., 2019], PubMed [Cohan et al., 2018], Reddit TIFU [Kim et al., 2019], SAMSum [Gliwa et al., 2019], and XSum [Narayan et al., 2018].
Peer-Review. ICLR reviews collected via OpenReview [ope, ]; see also the PeerRead corpus [Kang et al., 2018].
Agent Response Generation:
  • Model: GPT-4o-mini
  • Temperature: 0.7
  • Max tokens: 150 (summarization), 2000 (peer review), unbounded (translation)
  • Identical base prompts with condition-specific modifications
Mechanism Evaluation:
  • MI/GPPM: Llama 3.3-70B-Instruct for log probabilities
  • TVD-MI: GPT-4o-mini for categorical critic
  • Judge: GPT-4o-mini for pairwise comparison
  • All evaluations at temperature 0.7 for consistency
A.5 Computational Requirements
Our comprehensive evaluation involved:
  • 10 domains × 100-500 items × 30 conditions =  135,000 agent responses
  • 870 pairwise comparisons per item =  4.35 million evaluation calls
  • Approximately 500 million tokens processed
  • 72 hours of API computation time
Despite this scale, deployment requires only single API calls per evaluation, making our mechanisms practical for real-world use.
Appendix B Additional Findings
B.1 The Inverted-U Pattern: Compression and Information Structure
For compression ratio, quadratic models significantly outperformed linear fits for all primary mechanisms. GPPM showed the most improvement (R² increasing from 0.029 to 0.684, p = 0.007), while TVD-MI exhibited similar gains (R² from 0.046 to 0.674, p = 0.008). The quadratic coefficient was negative for all mechanisms, confirming the inverted-U shape with peaks at compression ratios of 9.6:1 (MI), 11.0:1 (GPPM), and 11.2:1 (TVD-MI).
B.2 LLM Judge Can Produce Inverted Evaluations
@@
Table 6 illustrates this inversion from specific examples from the CNN / Daily news summarization task. Human reference summaries, the gold standard for quality, received the lowest scores from LLM judges (0.234 with context, 0.117 without). Meanwhile, "Conspiracy Theory" summaries that fabricated information scored 0.703 with context and 0.777 without. This represents substantially higher scores for misinformation than for accurate content.
@@
Appendix C Peer Review Experiment Details
@@ C.1 Experimental Configuration
C.2 Agent Review Strategies
@@ 29 Checklist Review Standard checklist without deep engagement Low Effort
C.3 Two-Stage Review Generation Process
@@
C.3.1 Stage 1: Reading and Note-Taking